OpenStudy (anonymous):

square root 20ab divided by square root 45a^2b^3

OpenStudy (mertsj):

\[\frac{\sqrt{20ab}}{\sqrt{45a^2b^3}}=\sqrt{\frac{20ab}{45a^2b^3}}\]

OpenStudy (mertsj):

Think of it like that and simplify the fraction first